OpenStudy (anonymous):

Not sure how to do this: Determine algebraically whether the function is even, odd, or neither even nor odd. f(x) = -9 x^3 + 8x Would it be odd because the answer will always come out odd whether there is a -x or x?

OpenStudy (anonymous):

first off it is clearly "odd" because it is a polynomial with only odd exponents

OpenStudy (anonymous):

but your teacher wants you to do this: \[f(-x)=9(-x)^3+8(-x)=-9x^3-x=-(9x^3+x)=-f(x)\] and so since \[f(-x)=-f(x)\] it is odd

OpenStudy (anonymous):

Ok, got it thank you very much.

OpenStudy (anonymous):

typo there i meant \[f(-x)=9(-x)^3+8(-x)=-9x^3-8x=-(9x^3+8x)=-f(x)\]

OpenStudy (anonymous):

yw